OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

solve y = 3x - 6, so x = 2, and therefore y = 0. now you have two points (2, -4) and (2,0) as the equation is parallel to y = 3x - 6, this means the same gradient, consider it in the form y = mx + c, where m is the gradient so, m will equal 3 for the equation in question. Use y_2 - y_1 = m(x_2 - x_1) to determine it. 0 - (-4) = 3(2-2) ==> 4 = 0 which doesnt work, so i think you meant to put the point as (-2,-4) which would make it y = 3x - 12

OpenStudy (anonymous):

Ohhhh my bad! But i gt it now :) THANK YOU!!! :)) @Iamgmg90